In the short​ run, a tool manufacturer has a fixed amount of capital. Labor is a variable input. The cost and output structure that the firm faces is depicted in the following table.

Suppose that for the​ firm, the goods market is perfectly competitive. The market price of the product is ​$4 at each quantity supplied by the firm.  

Labor

Supplied

Total Physical

Product

Hourly Wage

Rate​ ($)

Total Wage

Cost

Marginal Factor

Cost

10

100

5

50

11

109

6

66

16

12

116

7

84

18

13

121

8

104

20

14

124

9

126

22

15

125

10

150

24

What is the amount of labor that this​ profit-maximizing firm will​ hire? __ workers. ​(Enter a numeric response using an​ integer.)
Answer #1
L TP MP= change in TP Marginal revenue product = (MR)(MP)= P(MP)= 4(MP) W Total wage cost Marginal factor cost
10 100 - 0 5 50 -
11 109 9 36 6 66 16
12 116 7 27 7 84 18
13 121 5 20 8 104 20
14 124 3 12 9 126 22
15 125 1 4 10 150 24

The profit maximization condition is where MRP = MFC.

So, the profit maximizing firm will hire 13 workers .